OpenStudy (sonic709):

Find f(-3) for f(x) = 3x - x² A. f(-3) = 18 B. f(-3) = -15 C. f(-3) = 0 D. f(-3) = -18

OpenStudy (amorfide):

if I had x+2=5 and I said that 5-2=x so 3=x then I said that f(x)=x+2 and find f(3) then you see from f(x) we have a function of x (the stuff in the brackets) and for f(3) we have a function of 3 so basically, let x=3 if I had f(4) let x=4 so in your case, f(-3), let x=-3, substitute it into the function, and that is your answer

OpenStudy (amorfide):

that is, f(-3)=3(-3) - (-3)^2